OpenStudy (anonymous):

What is the first value of of t greater than zero for which the derivative is greater than zero?

OpenStudy (anonymous):

\[\huge f'(x)\frac{ sintcost }{ 25 \pi }\]

OpenStudy (anonymous):

f'(x)= **

OpenStudy (anonymous):

Uh... there's no "first" value.

OpenStudy (anonymous):

that's what my question states :S

OpenStudy (anonymous):

Are they asking to solve for t, equation the derivative to zero?

OpenStudy (anonymous):

At t=0, the derivative will be 0. For 0<t<\(\frac{\Large\pi}{2}\) the derivative will be positive because sin and cos will both be positive.

OpenStudy (primeralph):

\[T = 0 + \left| \delta t \right|\]